identify the slope and y intercept
y=5/4x + 1

Answer by addingup(3677)

y = 5/4x + 1
The slope-intercept form is y = mx+b, where m is the slope and b is the y-intercept. Find the values of m and b using the form y = mx+b:
m = 5/4
b = 1
The slope of the line is the value of m, and the y-intercept is the value of b.
Slope: 5/4
Y-Intercept: 1